OpenStudy (anonymous):

PLEASE HELP WILL FAN AND GIVE MEDAL A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represent 60 miles? Enter your answer in the box. inches Part B How many miles are represented by 1 4 inch on the map? Enter your answer in the box. miles

OpenStudy (anonymous):

@amorfide can you help me please?

OpenStudy (anonymous):

i don't understand how to do it

OpenStudy (amorfide):

you know 2inches=25 miles how do you get from 25 to 60? 25 multiplied by an unknown is 60 so let's make the unknown x 25x=60 x=60/25 x=12/5 so use a calculator and you know x=2.4 so this means you multipled 25 by 2.4 to get to 60 so if you multiply 2inches by the same, then you get your amount in inches

OpenStudy (anonymous):

i don't understand so the answer is 36?

OpenStudy (anonymous):

@amorfide ?

OpenStudy (anonymous):

or is it the answer 2.4?

OpenStudy (anonymous):

u there @amorfide ?

OpenStudy (amorfide):

you know that 25 was multiplied by something to get to 60 so 25x=60 work out x so once you know x=2.4 25(2.4)=60 so to find out how many inches represent 60 miles you have 2x=25x 2(2.4)=25(2.4) so 4.8inches=60miles

OpenStudy (anonymous):

so the answer is 4.8?

OpenStudy (anonymous):

what about part B how do i do that? @amorfide

OpenStudy (amorfide):

yes 4.8 you wrote part b poorly I don't understand

OpenStudy (anonymous):